# Large Language Models and Code Security: A Systematic Literature Review

### v2

@@ VII-C Fine-tuning for Vulnerability Detection and Fixing
Several studies have focused on fine-tuning LLMs to enhance their performance in vulnerability detection and repair. These studies demonstrate that fine-tuning can improve the accuracy and effectiveness of LLMs in identifying and addressing software vulnerabilities.
@@
Mao et al. [106] proposed a framework named LLMVulExp, which utilizes fine-tuned LLMs trained to recognize, understand, and repair vulnerabilities, while offering detailed insights and repair suggestions. LLMVulExp can generate explanations of vulnerabilities with high accuracy, clarity, and actionability. The authors also explored using LLMs for automated data annotation to reduce manual effort. They found that. in certain scenarios, integrating vulnerability explanations into the fine-tuning process positively impacts performance in vulnerability detection. Last, they concluded that by directing LLMs to focus on key code snippets, the effectiveness of vulnerability detection can be significantly improved.
@@
Zhang et al. [107] fine-tuned LLMs on GitHub workflow tasks. They explored LLMs effectiveness in generating workflows, identifying defects, and repairing them, focusing on syntactic errors and code injection vulnerabilities. Although GPT-3.5 produced expected workflows, it occasionally generated defective ones. In contrast, Code Llama was less likely to generate expected workflows but had a higher probability of producing valid workflows. In terms of error detection, off-the-shelf StarChat achieved the best results in detecting syntactic errors, with a 100% F1 score, while fine-tuned GPT-3.5 performed best at detecting code injection vulnerabilities, with about 99% F1 score. However, LLMs can create workflows with code injection vulnerabilities, so it is important to be cautious when using workflows generated by LLMs.
@@
In light of what we have discussed so far, one work in the literature has looked at the issue of insecure LLM-generated code from a different angle. Storhaug et al. [108] proposed a novel approach aimed to reduce preemptively the amount of vulnerable code. Named vulnerability-constrained decoding, this technique involves using a small dataset of labeled vulnerable lines of code to fine-tune LLM, allowing it to include vulnerability labels when generating code. This approach effectively mitigates the risk of generating insecure code by constraining the model’s output during decoding.
VIII Prompting strategies for detecting and fixing vulnerabilities
@@
The effectiveness of LLMs for code vulnerability detection and fixing can vary significantly, depending on how the prompts are structured. Well-crafted prompts can guide LLMs towards more accurate and efficient solutions, while incomplete or ambiguous prompts may result in partial or ineffective solutions [46]. This section explores the impact of various prompting strategies on the detection and fixing performance of LLMs. To summarize the studies for the reader, Table VII lists the different prompting strategies, and corresponding studies.
@@
TABLE VII: Prompt engineering techniques used in the studies for vulnerability detection and fixing.
@@
Zero-shot Nong et al. [109], Steenhoek et al. [52], Pearce et al. [96], Mohajer et al. [44], Ullah et al. [55], Zhou et al. [66], Tamberg et al. [67], Zhang et al. [46], Nong et al. [98], Liu et al [110], Zhang et al. [107], Ni et al. [72], Liu et al [111], Mathews et al. [112]
Few-shot Nong et al. [109], Mohajer et al. [44], Ullah et al. [55], Zhou et al. [66], Liu et al.  [57], Tamberg et al. [67], Ding et al. [58], Zhang et al.  [46], Nong et al. [98], Zhang et al. [107], Gao et al. [63], Liu et al. [110]
Chain-of-Thought Yu et al. [54], Zhang et al. [68], Nong et al. [109], Steenhoek et al. [52], Zhou et al. [66], Liu et al. [57], Tamberg et al. [67], Ding et al. [58], Mao et al. [106], Nong et al [98], Li et al. [64], Yang et al. [113], Ni et al. [72], Kulsum et al. [114], Liu et al [110]
In-context Steenhoek et al. [52], Zhang et al. [46], Mathews et al. [112], Liu et al. [110], Li et al. [87], Ni et al. [72], Liu et al [111], Le et al. [8]
Task-oriented Ullah et al. [55], Zhou et al. [9], Bae et al. [115]
Role-oriented Ullah et al. [55], Zhou et al. [9], Çetin et al. [59], Zhang et al.  [46], Noever et al. [69], Pelofske et al. [60], Wen et al. [47], Fang et al. [90], Chen et al. [61], Yang et al. [113], Mohajer et al. [44], He et al. [11], Guo et al. [95], Ni et al. [72], Zhang et al. [68], Liu et al. [110]
VIII-A Zero-shot Prompting
Zero-shot prompting refers to the ability of an LLM to perform tasks without being provided with explicit task-specific examples. Instead of learning from examples, the model is simply given an instruction or a general guideline [116]. We provide an example of zero-shot prompting in Figure 4, demonstrating how the user gives instructions without providing any examples.
@@ VIII-A Zero-shot Prompting
Figure 4: Example of a prompt, using the zero-shot prompting technique. In green, the main body of the instructions. In blue, the code to be analyzed by the LLM.
Steenhoek et al. [52] experimented with three zero-shot approaches, but only one of them, according to our definition, can be classified as zero-shot. For example, one of the two other approaches, involved prompting the LLM to assess whether the given code contained vulnerabilities, with a fixed list of CWE types provided as references. We believe that this type of prompting does not align with zero-shot prompting, but in-context prompting (a technique further discussed in Subsection VIII-D). Their experiments demonstrated high performance in vulnerability detection with this type of prompt, achieving a balanced accuracy of about 62% when used on GPT-4.
@@
VIII-B Few-shot Prompting
While zero-shot prompting involves providing a model with only a general instruction or task description, few-shot prompting is a technique where the model is typically given a few task-specific examples, before performing the task [116], with n-shot prompting referring to cases that use a specific number n of examples. An example of a prompt using the few-shot technique is shown in Figure 5, where a set of task-specific examples are provided to the model, as part of the input.
@@ VIII-B Few-shot Prompting
Figure 5: Example of a prompt, using the few-shot prompting technique. In the gray boxes, the code examples and corresponding vulnerabilities provided to the LLM. In green, the main body of the instructions. In blue, the code to be analyzed.
In their research, [55] observed that few-shot prompting performs consistently better than zero-shot prompting, especially when tasked to detect vulnerabilities in code. In line with this, [66] noted that, among the different methods of utilizing ChatGPT, few-shot prompting is the most effective at detecting vulnerabilities. The intuition behind these consistent results, is that providing a model with explicit examples before performing a task, enhances its capability of recognizing patterns.
It is worth noting that other studies explored few-shot prompting [109, 44, 57, 98, 63, 72, 58, 110]. However, as we discuss in the following subsections, they did not achieve the same level of success, and deemed more promising other prompting techniques.
VIII-C Chain-of-Thought Prompting
Zhang et al. [68] explored CoT in a two-step process. First, they prompted ChatGPT to describe the functionality of a given code snippet. After the LLM provided an explanation, they prompted a follow-up question asking it to determine whether the code was vulnerable. The findings indicate that CoT prompting improves vulnerability detection accuracy for C/C++ by 21.6%. However, it results in a 4.6% decrease in accuracy for Java. Building on this, Nong et al. [109] instantiated CoT prompting through the Vulnerability-Semantics-guided Prompting (VSP) approach, which focuses on vulnerability semantics. VSP provides two main benefits. First, it isolates potentially vulnerable statements and their contextual relationships. Second, since LLMs (particularly transformer-based ones) struggle with extensive text, VSP narrows the scope and reduce unnecessary text. The results showed that VSP is effective across all tested models, significantly improving vulnerability detection accuracy. For instance, in the SARD dataset, VSP prompting enabled GPT-3.5 to achieve an accuracy of 97.65%, compared to the 65.88% obtained with standard prompting.
@@ VIII-C Chain-of-Thought Prompting
Example 1: <code1> - Takes user input as filename. - Opens the file specified by user input. - What could go wrong? If the user enters a malicious filename or path, it could point to sensitive files. - Conclusion: code is vulnerable to arbitrary file access. Now, analyze the following code for vulnerabilities: <code2>
@@
Figure 6: Example of a prompt, using the CoT prompting technique. In the gray boxes, the code example and the reasoning steps provided to the LLM. In green, the main body of the instructions. In blue, the code to be analyzed.
@@
Yang et al. [113] claimed that LLMs performance in identifying vulnerabilities is not satisfactory. So, in order to improve it, they propose a Deep Learning Augmented Prompting (DLAP) framework that combines DL and LLMs. Within the DLAP framework, various DL models were tested, and the authors reported that Linevul demonstrated the best performance. These DL models look at code and calculate probabilities of it containing certain errors. This information is combined into prompts that guide the LLM to focus on the right lines of code, making it implicitly fine-tuned (as defined by the authors). Additionally, the framework applies CoT prompting, guiding the LLM to think step-by-step and discard wrong paths. This dual approach improves the LLM problem-solving ability, making it more efficient and robust, minimizing at the same time performance drops and computation overhead. The researchers found that DLAP overall performance is superior to other prompting frameworks. Moreover, DLAP offers similar predictive performance to fine-tuning, but at a lower computational cost.
@@
Tamberg et al. [67] adopted CoT by developing a specialized eight-step CoT strategy, to mimic the manual checks performed by human beings during core reviews. The eight-step CoT strategy demonstrated optimal performance when paired with GPT-4, achieving an F1 score of 67%. Moreover, the authors experimented with using Tree of Thoughts (ToT) prompting, an advanced version of CoT. ToT relies on prompting an LLM to generate multiple possible responses at each step, then manually evaluating these responses, and selecting the best one [117]. Despite showing promise, ToT was deemed too computationally expensive, due to the large number of LLM calls required for generating and evaluating multiple candidates, at each reasoning step. Because of this limitation, it was only tested on a few CWEs using GPT-4 Turbo. For example, for CWE-23: Relative Path Traversal, ToT prompting achieved an F1 score of 75%. The authors suggested further exploration of ToT, to optimize its performance and address the computational challenges.
@@
Auxiliary information could also enhance CoT prompting. For instance, Yu et al. [54] found that integrating CWE references helps narrow down the problem domain, thereby improving the detection capabilities of LLMs. However, Bakhshandeh et al. [71] reported that including CWE information in prompts negatively affected GPT-3.5, reducing its accuracy in security defect detection. Similarly, Nong et al. [109] observed that GPT-3.5 often “forgets” the specific CWE it is tasked to identify, leading to incorrect reasoning during its analysis. On this ground, further research on this topic is warranted.
@@
Experimental findings from Liu et al. [57], indicate that CoT enhances vulnerability detection, while few-shot prompting negatively impacts it. Similarly, Mao et al. [106] proposed to incorporate CoT prompting within their framework for detecting and explaining vulnerabilities, showing that this approach can improve LLMs ability to identify and clarify security issues.
@@
Kulsum et al. [114] proposed VRpilot, an LLM-based vulnerability repair technique that uses CoT to reason about a vulnerability, before generating patch candidates. VRpilot iteratively adjust prompts, based on the error messages from the compiler and test suite. On several Java and C datasets, VRpilot produced 14% more correct patches for C and 7.6% more for Java, compared to the zero-shot vulnerability repair proposed by Pearce et al. [96].
@@
Dynamic adaptive prompting, proposed in [98], utilizes the CoT prompting strategy to incorporate CWE references into the prompts and generate code patches, accordingly. This approach improves LLMs performance in vulnerability patching, outperforming the same LLMs that use either standard prompting without guidance, or zero-shot prompting with CWE references. CoT prompting has also been studied in a few other papers [52, 66, 58, 64, 72, 110], but not as exhaustively as in the described studies.

### v3

@@ 7.3. Fine-tuning for Vulnerability Detection and Fixing
Several studies have focused on fine-tuning LLMs to enhance their performance in vulnerability detection and fixing. These studies demonstrate that fine-tuning can improve the accuracy and effectiveness of LLMs in identifying and addressing software vulnerabilities.
@@
Mao et al. (Mao et al., 2024) proposed LLMVulExp, a framework that uses fine-tuned LLMs to identify, explain, and fix vulnerabilities, while providing detailed insights and suggestions. LLMVulExp generates accurate, clear, and actionable explanations. The authors also explored automated data annotation with LLMs to reduce manual effort. They found that integrating vulnerability explanations into the fine-tuning process can improve detection performance, and guiding LLMs to focus on key code snippets further enhances effectiveness.
@@
Zhang et al. (Zhang et al., 2024c) fine-tuned LLMs on GitHub workflow tasks, evaluating their ability to generate workflows, identify defects, and apply fixes—focusing on syntactic errors and code injection vulnerabilities. Although GPT-3.5 produced expected workflows, it often generated defective ones. In contrast, Code Llama was less likely to generate expected workflows but had a higher probability of producing valid workflows. In terms of error detection, off-the-shelf StarChat achieved 100% F1 score when detecting syntax errors, while fine-tuned GPT-3.5 performed best at detecting code injection vulnerabilities, with about 99% F1 score. In conclusion, due to the potential presence of vulnerabilities, workflows generated by LLMs should be employed with caution.
@@
In light of what we have discussed so far, , one study approached the issue of insecure LLM-generated code from a different perspective. Storhaug et al. (Storhaug et al., 2023) proposed vulnerability-constrained decoding, a method that reduces the generation of vulnerable code by fine-tuning an LLM on a small dataset of labeled vulnerable lines. This allows the model to include vulnerability labels during generation, effectively constraining its output and mitigating security risks.
8. Prompting strategies for detecting and fixing vulnerabilities
@@
While LLMs show promise in detecting and fixing vulnerabilities, their effectiveness is not solely determined by model architecture or fine-tuning, it also heavily depends on how the prompts are structured. Well-crafted prompts can guide LLMs towards more accurate and efficient solutions, while incomplete or ambiguous prompts may result in partial or ineffective solutions (Zhang et al., 2024d). This section explores the impact of various prompting strategies on the detection and fixing performance of LLMs. To summarize the studies for the reader, Table 7 lists the different prompting strategies, and corresponding studies.
@@
Table 7. Prompt engineering techniques used in the studies for vulnerability detection and fixing.
@@
Zero-shot Nong et al. (Nong et al., 2024a), Steenhoek et al. (Steenhoek et al., 2024), Pearce et al. (Pearce et al., 2023), Mohajer et al. (Mohajer et al., 2024), Ullah et al. (Ullah et al., 2024), Zhou et al. (Zhou et al., 2024c), Tamberg et al. (Tamberg and Bahsi, 2024), Zhang et al. (Zhang et al., 2024d), Nong et al. (Nong et al., 2024b), Liu et al. (Liu et al., 2024b), Zhang et al. (Zhang et al., 2024c), Ni et al. (Ni et al., 2024), Liu et al. (Liu et al., 2024d), Mathews et al. (Mathews et al., 2024)
Few-shot Nong et al. (Nong et al., 2024a), Mohajer et al. (Mohajer et al., 2024), Ullah et al. (Ullah et al., 2024), Zhou et al. (Zhou et al., 2024c), Liu et al. (Liu et al., 2024a), Tamberg et al. (Tamberg and Bahsi, 2024), Ding et al. (Ding et al., 2025), Zhang et al.  (Zhang et al., 2024d), Nong et al. (Nong et al., 2024b), Zhang et al. (Zhang et al., 2024c), Gao et al. (Gao et al., 2023), Liu et al. (Liu et al., 2024b)
Chain-of-Thought Zhang et al. (Zhang et al., 2024b), Mohajer et al. (Mohajer et al., 2024), Steenhoek et al. (Steenhoek et al., 2024), Pearce et al. (Pearce et al., 2023), Mathews et al. (Mathews et al., 2024), Nong et al. (Nong et al., 2024a), Yu et al. (Yu et al., 2024), Ullah et al. (Ullah et al., 2024), Zhou et al. (Zhou et al., 2024c), Tamberg et al. (Tamberg and Bahsi, 2024), Zhang et al. (Zhang et al., 2024d), Nong et al. (Nong et al., 2024b), Liu et al. (Liu et al., 2024b), Zhang et al. (Zhang et al., 2024c), Ni et al. (Ni et al., 2024), Liu et al. (Liu et al., 2024d)
In-context Steenhoek et al. (Steenhoek et al., 2024), Zhang et al. (Zhang et al., 2024d), Mathews et al. (Mathews et al., 2024), Liu et al. (Liu et al., 2024b), Li et al. (Li et al., 2024a), Ni et al. (Ni et al., 2024), Liu et al (Liu et al., 2024d), Le et al. (Le et al., 2024)
Task-oriented Ullah et al. (Ullah et al., 2024), Zhou et al. (Zhou et al., 2024d), Bae et al. (Bae et al., 2024)
Role-oriented Ullah et al. (Ullah et al., 2024), Zhou et al. (Zhou et al., 2024d), Çetin et al. (Çetin et al., 2024), Zhang et al. (Zhang et al., 2024d), Noever et al. (Noever, 2023), Pelofske et al. (Pelofske et al., 2024), Fang et al. (Fang et al., 2024), Chen et al. (Chen et al., 2024a), Yang et al. (Yang et al., 2025), Mohajer et al. (Mohajer et al., 2024), He et al. (He and Vechev, 2023), Guo et al. (Guo et al., 2024), Ni et al. (Ni et al., 2024), Zhang et al. (Zhang et al., 2024b), Liu et al. (Liu et al., 2024b), Wen et al. (Wen et al., 2024a)
8.1. Zero-shot Prompting
Zero-shot prompting refers to the ability of an LLM to perform tasks without being provided with explicit task-specific examples. Instead of learning from examples, the model is simply given an instruction or a general guideline (Kojima et al., 2022). We provide an example of zero-shot prompting in Figure 4, demonstrating how the user gives instructions without providing any examples.
@@ 8.1. Zero-shot Prompting
Figure 4. Example of a prompt, using the zero-shot prompting technique. In green, the main body of the instructions. In blue, the code to be analyzed by the LLM.
Steenhoek et al. (Steenhoek et al., 2024) tested three zero-shot approaches, but only one of them, according to our definition, can be classified as zero-shot. For example, one of the two other approaches, involved prompting the LLM to assess whether the given code contained vulnerabilities, with a fixed list of CWE types provided as references. We believe that this type of prompting does not align with zero-shot prompting, but in-context prompting (a technique further discussed in Subsection 8.4). Their experiments demonstrated high performance in vulnerability detection with this type of prompt, achieving a balanced accuracy of about 62% when used on GPT-4.
@@
8.2. Few-shot Prompting
While zero-shot prompting involves providing a model with only a general instruction or task description, few-shot prompting is a technique where the model is typically given a few task-specific examples, before performing the task (Kojima et al., 2022), with n-shot prompting referring to cases that use a specific number n of examples. An example of a prompt using the few-shot technique is shown in Figure 6, where a set of task-specific examples are provided to the model, as part of the input.
In their research, (Ullah et al., 2024) observed that few-shot prompting performs consistently better than zero-shot prompting, especially when tasked to detect vulnerabilities in code.
In line with this, (Zhou et al., 2024c) noted that, among the different methods of utilizing ChatGPT, few-shot prompting is the most effective at detecting vulnerabilities. The intuition behind these consistent results, is that providing a model with explicit examples before performing a task, enhances its capability of recognizing patterns.
It is worth noting that other studies explored few-shot prompting (Nong et al., 2024a; Mohajer et al., 2024; Liu et al., 2024a; Nong et al., 2024b; Gao et al., 2023; Ni et al., 2024; Ding et al., 2025; Liu et al., 2024b). However, as we discuss in the following subsections, they did not achieve the same level of success, and deemed more promising other prompting techniques.
8.3. Chain-of-Thought Prompting
@@
Figure 5. Example of a prompt, using the few-shot prompting technique. In the gray boxes, the code examples and corresponding vulnerabilities provided to the LLM. In green, the main body of the instructions. In blue, the code to be analyzed.
@@
Example 1: <code1> - Takes user input as filename. - Opens the file specified by user input. - What could go wrong? If the user enters a malicious filename or path, it could point to sensitive files. Conclusion: Code is vulnerable to arbitrary file access. Now, analyze the following code for vulnerabilities: <code2>
@@
Figure 6. Example of a prompt, using the CoT prompting technique. In the gray boxes, the code example and the reasoning steps provided to the LLM. In green, the main body of the instructions. In blue, the code to be analyzed.
Zhang et al. (Zhang et al., 2024b) explored CoT in a two-step process. First, they prompted ChatGPT to describe the functionality of a given code snippet. After the LLM provided an explanation, they prompted a follow-up question asking it to determine whether the code was vulnerable. The findings indicate that CoT prompting improves vulnerability detection accuracy for C/C++ by 21.6%. However, it results in a 4.6% decrease in accuracy for Java. Building on this, Nong et al. (Nong et al., 2024a) instantiated CoT prompting through the Vulnerability-Semantics-guided Prompting (VSP) approach, which focuses on vulnerability semantics. VSP provides two main benefits. First, it isolates potentially vulnerable statements and their contextual relationships. Second, since LLMs (particularly transformer-based ones) struggle with extensive text, VSP narrows the scope and reduce unnecessary text. The results showed that VSP is effective across all tested models, significantly improving vulnerability detection accuracy. For instance, in the SARD dataset, VSP prompting enabled GPT-3.5 to achieve an accuracy of 97.65%, compared to the 65.88% obtained with standard prompting.
@@
Yang et al. (Yang et al., 2025) claimed that LLMs performance in identifying vulnerabilities is not satisfactory. So, in order to improve it, they propose a Deep Learning Augmented Prompting (DLAP) framework that combines DL and LLMs. Within the DLAP framework, various DL models were tested, and the authors reported that Linevul demonstrated the best performance. These DL models look at code and calculate probabilities of it containing certain errors. This information is combined into prompts that guide the LLM to focus on the right lines of code, making it implicitly fine-tuned (as defined by the authors). Additionally, the framework applies CoT prompting, guiding the LLM to think step-by-step and discard wrong paths. This dual approach improves the LLM problem-solving ability, making it more efficient and robust, minimizing at the same time performance drops and computation overhead. The researchers found that DLAP overall performance is superior to other prompting frameworks. Moreover, DLAP offers similar predictive performance to fine-tuning, but at a lower computational cost.
@@
Tamberg et al. (Tamberg and Bahsi, 2024) adopted CoT by developing a specialized eight-step CoT strategy, to mimic the manual checks performed by human beings during core reviews. The eight-step CoT strategy demonstrated optimal performance when paired with GPT-4, achieving an F1 score of 67%. Moreover, the authors experimented with using Tree of Thoughts (ToT) prompting, an advanced version of CoT. ToT relies on prompting an LLM to generate multiple possible responses at each step, then manually evaluating these responses, and selecting the best one (Yao et al., 2024b). Despite showing promise, ToT was deemed too computationally expensive, due to the large number of LLM calls required for generating and evaluating multiple candidates, at each reasoning step. Because of this limitation, it was only tested on a few CWEs using GPT-4 Turbo. For example, for CWE-23: Relative Path Traversal, ToT prompting achieved an F1 score of 75%. The authors suggested further exploration of ToT, to optimize its performance and address the computational challenges.
@@
Auxiliary information could also enhance CoT prompting. For instance, Yu et al. (Yu et al., 2024) found that integrating CWE references helps narrow down the problem domain, thereby improving the detection capabilities of LLMs. However, Bakhshandeh et al. (Bakhshandeh et al., 2023) reported that including CWE information in prompts negatively affected GPT-3.5, reducing its accuracy in security defect detection. Similarly, Nong et al. (Nong et al., 2024a) observed that GPT-3.5 often “forgets” the specific CWE it is tasked to identify, leading to incorrect reasoning during its analysis. On this ground, further research on this topic is warranted.
@@
Experimental findings from Liu et al. (Liu et al., 2024a), indicate that CoT enhances vulnerability detection, while few-shot prompting negatively impacts it. Mao et al. (Mao et al., 2024) proposed to incorporate CoT prompting within their framework for detecting and explaining vulnerabilities, showing that this approach can improve LLMs ability to identify and clarify security issues.
@@
Kulsum et al. (Kulsum et al., 2024) proposed VRpilot, an LLM-based vulnerability repair technique that uses CoT to reason about a vulnerability, before generating patch candidates. VRpilot iteratively adjust prompts, based on the error messages from the compiler and test suite. On several Java and C datasets, VRpilot produced 14% more correct fixes for C and 7.6% more for Java, compared to the zero-shot vulnerability repair proposed by Pearce et al. (Pearce et al., 2023).
@@
Dynamic adaptive prompting, proposed in (Nong et al., 2024b), utilizes the CoT prompting strategy to incorporate CWE references into the prompts and generate code patches, accordingly. This approach improves LLMs performance in vulnerability fixing, outperforming the same LLMs that use either standard prompting without guidance, or zero-shot prompting with CWE references. CoT prompting has also been studied in a few other papers (Steenhoek et al., 2024; Zhou et al., 2024c; Ding et al., 2025; Li et al., 2023; Ni et al., 2024; Liu et al., 2024b), but not as exhaustively as in the described studies.